OpenStudy (anonymous):

Need help!!! The perimeter of a basketball court is 84 meters and the length is 6 meters longer than twice the width.. what are the length and width

OpenStudy (anonymous):

L=2W+6 P=2L+2W 84=2(2W+6)+2W 84=4W+12+2W 84=6W+12 6W=84-12 6W=72 W=72/6 W=12 ans. for the width. L=2*12+6=24+6=30 ans. for the length. Proof: 84=2*30+2*12 84=60+24 84=84

OpenStudy (anonymous):

Hope it helps.

OpenStudy (anonymous):

12 is the width, 30 is the length.

OpenStudy (anonymous):

Umm can I get a reply saying it helped or its incorrect or something?